OpenStudy (anonymous):

limit x->0, tan2x/x find the limit, help!!

OpenStudy (anonymous):

l'Hopital's rule top and bottom. (2sec(x)^2)/1, limx->0 of 2sec(x)^2

OpenStudy (anonymous):

we havent learned that rule, yet, so u wouldnt have tan=cos/sin?

OpenStudy (anonymous):

sin/cos*

OpenStudy (turingtest):

is it\[\tan(2x)\]or\[\tan^2(x)\]?

OpenStudy (anonymous):

tan2x not ^2

OpenStudy (anonymous):

I wonder, I don't think it's possible to solve tan(2x)/x without l'Hopital's. Unless I'm derping. Maybe some weird half angle identity?

OpenStudy (anonymous):

my calc class just started last week n my professor has mentioned that rule but has yet to teach it, so wa does that rule state?

OpenStudy (anonymous):

http://mathworld.wolfram.com/LHospitalsRule.html Apparently I spelled it wrong, but w/e. The proof is lovely.

OpenStudy (anonymous):

you could try \[\lim_{x\to 0}\frac{\sin(2x)}{x\cos(2x)}\]

OpenStudy (anonymous):

thats as far as i got what do i after that satel.?

OpenStudy (anonymous):

break into to parts \[\lim_{x\to 0}\frac{1}{\cos(2x)}= 1\] and \[\lim_{x\to 0}\frac{\sin(2x)}{x}=2\]

OpenStudy (turingtest):

How about\[\frac{\tan(2x)}x=\frac{\sin(2x)}{x\cos(2x)}=2\frac{\sin x}{x}\cdot\frac{\cos x}{2\cos^2x-1}=2\]sat beat me as usual...

OpenStudy (anonymous):

yeah but you are quicker because i did not explain my second limit

OpenStudy (anonymous):

Man, algebra. I forget that if I can't solve a problem, it's probably just me doing algebra badly.

OpenStudy (zarkon):

I would write \[\frac{\sin(2x)}{x}=2\frac{\sin(2x)}{2x}\]

OpenStudy (turingtest):

dang...lol

OpenStudy (anonymous):

or replace x by x/2

OpenStudy (anonymous):

turning test, so 2cos^2x-1 cancels the top, n ur left wit 2?

OpenStudy (turingtest):

I didn't even need to do that identity, don't know why I did but you can just plug in x=0 to that second part and get 1

OpenStudy (anonymous):

the answer is 2 tho

OpenStudy (turingtest):

that's what you get in total\[\frac{\tan(2x)}x=\frac{\sin(2x)}{x\cos(2x)}=2\frac{\sin x}{x}\cdot\frac{\cos x}{\cos(2x)}=2\cdot1\cdot1=2\]

OpenStudy (turingtest):

I didn't write the limits in ...lazy I know

OpenStudy (anonymous):

got it!!